VEGAS Team Meeting

Attendees: Amanda, Joe, Marty, Richard, Mark, Randy, Dan, John, Hong, Toney

Agenda:

  1. DONE Rebuild of BOFs for specific frequencies; associated updates to vegas.configurations
  2. Mode 10--13,15--18 Testing
  3. DONE VEGAS Data Display
  4. Mode 10-19 (ghost frequencies issue)
  5. Blanking of last switched integration
  6. Status of “Manager writes to many sub-bands in 8 subband mode” problem (release and test)
  7. Content of August 'release'
  8. Blanking times
  9. ADC calibration routines and documentation for DIBAS
  10. Use of test time (including overnights)
  11. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Meeting notes:

  1. Rebuild of BOFs for specific frequencies; associated updates to Vegas configurations
    • If we build bofs for correct sample frequency then ADC issues (MMCM failures) should be resolved
    • In process - could test tomorrow
      • Hong will build (4.5 hrs per)
      • John will send needed clock frequencies with priorities
    • Changes to yellow block - a new library should be used - Hong will circulate library to be used for subsequent compiles.
      • has most recent ADC developments
      • Should be an easy build
  2. Mode 10--13,15--18 Testing
    • Testing was successful
  3. VEGAS Data Display
    • Gathering science input in meeting this week to inform work by Joe Masters
    • Add Jay. Maybe Sam, & Adam
  4. Mode 10-19 (ghost frequencies issue)
    • We have a workaround for the problems for modes 10-13 and 15-18
    • 14 and 19 modes are only one ones left to be fixed Once they are fixed the workaround above will be replaced
    • L1 and L8 problems identified by Randy
      • data ordering problem in L1.
      • works in simulation - compiled versions don't work (no data output)
  5. Blanking of last switched integration
    • Fixed and tested
  6. Status of “Manager writes to many sub-bands in 8 subband mode” problem (release and test)
    • successful
  7. Content of August 'release' for next semester observing
    • Currently: everything except Mode 14 & 19 based upon future testing
  8. Blanking times
    • Need to set up maximum for switching times to optimize observational efficiencies
    • Noise diode
      • currently 2ms for noise diode
      • use values from 'Joe's document' for published value
    • LO - to be validated/updated by Joe
      • Frequency switching w/o doppler tracking should be 8ms (update MR 2Q32014)
      • Frequency switching w/doppler tracking should be 80ms (update MR 2Q32014)
  9. ADC calibration routines and documentation for DIBAS
    • Could be cleared by #1 above
  10. Use of test time (including overnights)
There is a possibility of a moving telescope Wednesday night and Thursday daytime – to be confirmed.
    • Unscheduled request(s)
      • none
    • Tuesday 08:00 - 13:30
    • Wednesday All day
      • 12:00 - 16:00 - bof testing [Prestage]
      • Wednesday eve. - on-sky testing - mixed mode testing needs minimum of 10 minutes data per permutation [Prestage]
    • Thursday All day [pending]
      • 08:00 - 12:00 - servo tests [Prestage]
    • Monday overnight
      • UTC delta tests? [Roshi]
  1. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-07-01
Topic revision: r3 - 2014-07-01, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback